# CoPRS: Learning Positional Prior from Chain-of-Thought for Reasoning Segmentation

### v1

@@ 4.4 Ablation Study (RQ3)
To gain a deeper understanding of the contributing factors, we run ablations with CoPRS-7B on RefCOCO+, RefCOCOg and ReasonSeg. We examine training modes, GRPO hyperparameters, and segmentation loss configurations, and report their impact on performance. These studies provide insight into the effectiveness and stability of our method.
@@
Training Modes. We compare reinforcement learning, segmentation supervision, and a combined objective for CoPRS-7B. As shown in Section 4.4, the combined objective achieves the best performance. This suggests that reinforcement learning strengthens reasoning, while supervised signals sharpen mask generation. Together they are more effective for complex reasoning segmentation.
GRPO Training Hyperparameters. Next, we study the effects of sampling number and reward design in GRPO training. Sampling number is the number of responses sampled per question during rollout. As shown in Section 4.4, increasing sampling number improves performance on both in domain (RefCOCO+, RefCOCOg) and out of distribution (ReasonSeg) data. This trend is expected, as larger groups broaden exploration, yielding more stable updates. We also evaluate alternative reward designs. Section 4.4 compares combinations of the format score and the segmentation score. Compared with the format score, the segmentation score contributes more to segmentation performance. Concurrently, an appropriate format score improves generalization on out of distribution data.
@@
Segmentation loss combinations. We compare segmentation loss configurations with varying coefficients Section 4.4 to assess the contribution of each component, while keeping the BCE weight fixed at 1. Adding a focal term, which emphasizes hard pixels and fine-grained structures, improves segmentation performance. However, the ratio of weight coefficients between focal loss and dice loss needs to be tuned to improve both global and local mask quality.
Table 3: Effect of training mode.
Reinforcement
Learning
Segmentation
Supervision
RefCOCO+ RefCOCOg ReasonSeg
67.5 55.3 32.9
77.1 64.8 50.7
80.3 67.2 55.1
Table 4: Effect of sampling number.
Sampling
Number
RefCOCO+ RefCOCOg ReasonSeg
2 76.5 64.2 51.0
4 78.3 66.8 53.0
8 80.3 67.2 55.1
Table 5: Effect of reward function.
Format
Score
Segmentation
Score
RefCOCO+ RefCOCOg ReasonSeg
0.0 1.0 80.1 66.8 53.4
1.0 0.0 77.0 65.1 51.0
0.5 0.5 79.9 67.0 54.9
0.3 0.7 80.3 67.2 55.1
Table 6: Effects of segmentation losses.
Dice
Loss
Focal
Loss
RefCOCO+ RefCOCOg ReasonSeg
1.5 0 77.9 64.1 53.0
1.5 10 79.2 65.1 54.2
3.0 20 79.8 65.2 54.8
3.0 10 80.3 67.2 55.1
@@ 5 Conclusions
In this work, we propose CoPRS, connecting language reasoning with segmentation via an interpretable and differentiable interface. CoPRS implements this idea with a learnable concentration query that produces the heatmap as a positional prior, from which precise masks are decoded, within a unified training framework combining reinforcement learning and segmentation supervision. This interface avoids feeding hidden features to the decoder or representing positions in text, instead providing a direct, interpretable link between MCoT and mask generation. Empirically, CoPRS matches or surpasses the best reported metrics on multiple datasets. Further analysis shows that heatmap quality strongly correlates with final mask accuracy and sample visualizations show the same pattern. Overall, CoPRS delivers strong concentration from reasoning and high mask quality in a clear and unified formulation, and provides a clean starting point for future work on perception aligned with instructions.
@@
Reproducibility Statement. We point readers to the fundamental setup in Experimental Setup (Section 4.1) and to the appendix Implementation Details (Appendix B), which concisely summarizes the pipeline implementation (Section B.1) and the training configuration (Section B.2). These sections contain the information needed to reproduce our results. We will release code, configurations, and checkpoints upon acceptance.
@@ A.1 Group Relative Policy Optimization
where $\pi_{\bm{\theta}}$ and $\pi_{{\bm{\theta}}_{old}}$ are the current and old policy models, and $q,o$ are questions and outputs sampled from the question dataset and the old policy $\pi_{{\bm{\theta}}_{old}}$, respectively. $\varepsilon$ is a clipping-related hyper-parameter introduced in PPO for stabilizing training. The advantage, $A_{t}$, is based on the reward $\{r_{\geq t}\}$ and a learned value function $V_{\psi}$, computed by applying Generalized Advantage Estimation (GAE) (Schulman et al., 2015). Furthermore, a per-token KL penalty from a reference model is added in the reward at each token to mitigate over-optimization of the reward model (Ouyang et al., 2022), denoted as:
@@
where $r_{\varphi}$ is the reward model, $\pi_{\bm{\theta}}$ is the reference model, which is usually the initial policy model, and $\beta$ is the coefficient of the KL penalty.
PPO relies on a separate value function that is typically another model of comparable size as the policy model, imposing heavy memory and computational costs. Additionally, the value function is treated as a baseline in the calculation of the advantage for variance reduction. Moreover, in the LLM context, usually only the last token is assigned a reward score by the reward model, which may complicate the training of a value function that is accurate at each token. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is proposed to address these drawbacks via obviating the need for additional value function approximation as in PPO, and using the average reward of multiple sampled outputs, produced in response to the same question, as the baseline. Specifically, for each question $q$, GRPO samples a group of outputs $\{o_{1},o_{2},\cdots,o_{G}\}$ from the old policy $\pi_{{\bm{\theta}}_{old}}$ and then optimizes the policy model by maximizing the following objective:
@@
where $\varepsilon$ and $\beta$ are hyper-parameters, and $\hat{A}_{i,t}$ is the advantage calculated based on relative rewards of the outputs inside each group only. For each question $q$, a group of outputs $\{o_{1},o_{2},\cdots,o_{G}\}$ are sampled from the old policy model $\pi_{{\bm{\theta}}_{old}}$. The score of the outputs is obtained through a reward model, yielding $G$ rewards $\{r_{1},r_{2},\cdots,r_{G}\}$ correspondingly. The advantages  $\hat{A}_{i,t}$ of all tokens in the output is the normalized reward, i.e., $\hat{A}_{i,t}=\tilde{r}_{i}=\frac{r_{i}-\text{mean}(\bm{r})}{\text{std}(\bm{r})}$. In addition, GRPO directly adds the KL divergence between the trained policy and the reference policy to the loss, avoiding complicating the calculation of $\hat{A}_{i,t}$. The KL divergence is estimated by the following unbiased estimator:
@@ A.2 Additional Related Work
Multimodal Chain-of-Thought. Multimodal chain-of-thought (MCoT) (Wang et al., 2025) reasoning has recently attracted substantial attention, particularly in its integration with MLLMs. Early implementations, such as Multimodal-CoT (Zhang et al., 2024c), has established a basic MCoT pattern by generating intermediate rationales before predictions. MC-CoT (Tan et al., 2024) further refines this paradigm by employing word-level majority during training to enhance the quality of generated rationals. The dependence on high-quality MCoT training data hinders the further improvement of the inference ability of traditional methods. Most recently, the great success of Deepseek-R1 (Guo et al., 2025) has provided a way (i.e. GRPO) to enhance LLM inference capabilities through model autonomous exploration without the need for expensive CoT annotation data. Inspired by this, subsequent works utilize the GRPO strategy to efficiently enhance the reasoning ability of MLLMs. For example, Vision-R1 (Huang et al., 2025) first utilizes existing MLLM and DeepSeek-R1, as well as data filtering, through modal bridging to generate multimodal cold start CoT data, and then applies GRPO to further enhance the model’s inference capability. Perception-R1  (Yu et al., 2025a) explores the effects of RL on different perception tasks and optimizes the reward modeling to support perception policy learning. In addition, Chain-of-Shot (Hu et al., 2025) further extends GRPO strategy to optimize frame sampling via binary video summaries. In this work, we study a heatmap-based positional prior that couples MCoT with precise positional perception in a unified training framework for GRPO strategy and segmentation supervision, addressing the gap between high-level reasoning and pixel-level segmentation.
@@ B.1 Pipeline Implementation
FSDP Workers. We precompute image features offline to reduce compute, so the frozen vision backbone is excluded from the training loop. Our framework uses three FSDP workers. (i): The actor contains all trainable modules (the MLLM and the segmentation components) and is responsible for parameter updates. (ii): The rollout worker runs the MLLM only, taking image and text inputs to generate responses via next token prediction. (iii): The frozen reference worker runs an MLLM as the reference policy to compute the KL term in $\mathcal{L}_{\text{{grpo}}}$ (eq. 7) and includes the segmentation modules to decode masks used for computation of mask reward scores and group advantages.
@@
B.2 Training configuration
@@
B.3 LLM usage statement
@@ Appendix C Appendix: Additional Sample Visualizations
Successful Cases. In Figure 6, we present Instances from the same category and those relevant to the instruction show elevated response in the heatmap (yellow regions). More importantly, the heatmap concentrates on the instance specified by the instruction, producing a sharp peak over the target (deep red regions). This concentration guides the decoder, yielding masks with accurate boundaries. TThese results indicate that the MLLM reasons over the image and text input, identifies the correct referent, while the positional prior concentrates the instances for futher precise mask prediction.
@@
In Figure 5, the first two rows depict scenes with many nearby instances, while the last three rows contain very small targets. Two failure modes emerge. (i) Resolution bottleneck: the positional prior is computed at 256×256 and the SAM embeddings at 64×64; when the longer image side exceeds 2k pixels, tiny objects can vanish after resizing and the decoder cannot reliably recover them. (ii) Same class crowd ambiguity: Same class crowd ambiguity: in dense groups of similar objects (e.g., crowds of people), the positional prior often spreads across many candidates with weak contrast, suggesting that a text only instruction is insufficient to disambiguate near duplicates and that the model has not fully learned the subtle semantic cues needed. These observations suggest that higher resolution inputs or multi-scale features, together with stronger instance level language grounding, are likely to improve performance on such cases.

### v2

@@ 4.4 Ablation Study (RQ3)
To gain a deeper understanding of the contributing factors, we perform ablation studies on RefCOCO+ with different MLLM backbones and varied vision backbones, and further ablations of CoPRS-7B on RefCOCO+, RefCOCOg, and ReasonSeg. We systematically examine MLLM backbone choice, vision backbone choice, GRPO group size, training mode, reward coefficients, and segmentation loss combinations.
Table 4: Effect of MLLM Backbone Choice. Gray row denotes the default backbone.
Method Backbone val testA testB
CoPRS-3B Qwen2.5-VL 71.8 78.9 66.5
CoPRS-7B Qwen2.5-VL 75.9 80.3 69.7
CoPRS-7B LLaVA-1.5 73.1 79.0 66.4
CoPRS-13B LLaVA-1.5 75.5 80.3 70.7
MLLM Backbone. For ablating the MLLM backbone, we additionally train CoPRS with LLaVA-1.5-7B/13B on RefCOCO+. Table 4 reports cIoU metrics of CoPRS versions with both LLaVA-1.5 and Qwen2.5-VL series. As expected, performance increases with backbone capacity, but the gains across different MLLM backbones are relatively modest. This indicates that CoPRS is not sensitive to the specific MLLM architecture and that our improvements largely transfer across different backbone choices. Together with the comparisons to prior work under the same LLaVA-1.5 backbone (Table 1), this suggests that our gains are complementary to backbone strength rather than being tied to a particular MLLM.
Table 5: Effect of Vision Backbone Choice. Gray row denotes the default backbone.
Backbone #Params(B) val testA testB
ViT-B 8.38 73.2 77.3 67.0
ViT-L 8.60 74.8 78.9 68.5
ViT-H 8.93 75.9 80.3 69.7
Vision Backbone. As shown in Table 5, we ablate SAM backbones (ViT-B/L/H) on RefCOCO+ with a fixed Qwen2.5-VL-7B MLLM and report the total parameters of the full pipeline. Larger vision backbones bring slightly better segmentation performance, but the improvement is modest and the overall trend remains stable across sizes. Additionally, vision backbones constitute only a small portion of the total parameters, so scaling them up only marginally increases overall computational cost.
(a) Effect of GRPO group size
(b) Effect of training mode choice
(c) Effect of mask reward coefficient.
(d) Effect of segmentation loss coefficients.
Figure 6: Ablation studies on GRPO group size, training mode, mask reward coefficient, and segmentation loss coefficients. (a) is evaluated on all splits of RefCOCO+, while (b)–(d) are evaluated on the test split of each dataset. Bold x-axis labels mark the default settings.
GRPO Group Size. We study the effects of GRPO group size during training. The group size $G$ denotes the number of responses sampled per question during rollout. As shown in Figure 6(a), increasing $G$ improves performance across splits of RefCOCO+. To quantify efficiency, we also report the total number of GRPO samples required to reach convergence (loss fluctuation $10\%$ over 300 steps) for $G\in\{2,4,8,16\}$. Particularly, the number of samples for convergence does not grow linearly with $G$, because larger groups offer more diverse candidates per step, improving exploration and the contrast between positive and negative samples. Empirically, we find that $G=8$ strikes a good trade-off between efficiency and performance.
@@
Training Modes. We compare reinforcement learning, segmentation supervision, and a combined objective for CoPRS-7B. As shown in Figure 6(b), the combined objective achieves the best performance. This suggests that reinforcement learning strengthens reasoning, while supervised signals sharpen mask generation. Together they are more effective for complex reasoning segmentation.
Reward Coefficients. We evaluate the impact of reward mixing ratio between mask reward score and format score. Figure 6(c) compares their combinations, where the format score is one minus the mask score. As the coefficient on the mask reward increases from 0 to 0.7, cIoU improves across all three datasets, but pushing it further to 1.0 slightly degrades performance. This pattern suggests that the segmentation term is the main driver of segmentation quality, while keeping a small contribution from the format score helps regularize the policy and improves generalization, especially on out of distribution data (ReasonSeg). We set the 0.7/0.3 weighting by default, with the segmentation reward dominant and the format score acting as a regularizer, and Figure 6(c) supports this choice.
@@
Segmentation Loss Combinations. We compare segmentation loss configurations with varying coefficients (see Figure 6(d)) to assess the contribution of each component, with BCE weight fixed at 1. To avoid the prohibitive cost of LLM experiments, we only probe a few representative weight settings, which already show trends consistent with our expectations. Adding a focal loss term, which emphasizes hard pixels and fine-grained structures, improves segmentation performance. The relative weight between focal and dice loss also affects the balance between global and local mask quality.
@@ 5 Conclusions
In this work, we propose CoPRS, connecting language reasoning with segmentation via an interpretable and differentiable interface. CoPRS implements this idea with a learnable concentration query to produce a positional prior instantiated as a heatmap, from which precise masks are decoded, within a unified framework combining reinforcement learning and segmentation supervision. This interface avoids feeding hidden features to the decoder or representing positions in text, instead providing a direct, interpretable alignment between reasoning and mask generation. Empirically, CoPRS attains strong performance across datasets. Further analysis shows that CoT trajectory and heatmap quality strongly correlate with final mask accuracy, and sample visualizations show the same pattern. Overall, CoPRS delivers strong concentration from reasoning and predicts precise masks in a unified formulation, providing a starting point for perception aligned with instructions.
@@
Reproducibility Statement. We point readers to the fundamental setup in Experimental Setup (Section 4.1), and to the appendix Implementation Details (Appendix B), which concisely summarizes the pipeline implementation (Section B.1), the design details (Section B.2) and the training configuration (Section B.3). These sections contain the information needed to reproduce our results. We will release code, configurations, and checkpoints upon acceptance.
@@ A.1 Group Relative Policy Optimization
where $\pi_{\bm{\theta}}$ and $\pi_{{\bm{\theta}}_{old}}$ are the current and old policy models, and $q,o$ are questions and outputs sampled from the question dataset and the old policy $\pi_{{\bm{\theta}}_{old}}$, respectively. $\varepsilon$ is a clipping-related hyper-parameter introduced in PPO for stabilizing training. The advantage, $A_{t}$, is based on the reward $\{r_{\geq t}\}$ and a learned value function $V_{\psi}$, computed by applying Generalized Advantage Estimation (GAE) (Schulman et al., 2015). Furthermore, a per-token KL penalty from a reference model is added to the reward at each token to mitigate over-optimization of the reward model (Ouyang et al., 2022), denoted as:
@@
where $r_{\varphi}$ is the reward model, $\pi_{\text{ref}}$ is the reference model, which is usually the initial policy model, and $\beta$ is the coefficient of the KL penalty.
PPO relies on a separate value function that is typically another model of comparable size to the policy model, imposing heavy memory and computational costs. Additionally, the value function is treated as a baseline in the calculation of the advantage for variance reduction. Moreover, in the LLM context, usually only the last token is assigned a reward score by the reward model, which may complicate the training of a value function that is accurate at each token. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is proposed to address these drawbacks by obviating the need for additional value function approximation as in PPO, and using the average reward of multiple sampled outputs, produced in response to the same question, as the baseline. Specifically, for each question $q$, GRPO samples a group of outputs $\{o_{1},o_{2},\cdots,o_{G}\}$ from the old policy $\pi_{{\bm{\theta}}_{old}}$ and then optimizes the policy model by maximizing the following objective:
@@
where $\varepsilon$ and $\beta$ are hyper-parameters, and $\hat{A}_{i,t}$ is the advantage calculated based on relative rewards of the outputs inside each group only. For each question $q$, a group of outputs $\{o_{1},o_{2},\cdots,o_{G}\}$ are sampled from the old policy model $\pi_{{\bm{\theta}}_{old}}$. The score of the outputs is obtained through a reward model, yielding $G$ rewards $\{r_{1},r_{2},\cdots,r_{G}\}$ correspondingly. The advantages  $\hat{A}_{i,t}$ for all tokens in an output are defined as the normalized reward, i.e., $\hat{A}_{i,t}=\tilde{r}_{i}=\frac{r_{i}-\text{mean}(\bm{r})}{\text{std}(\bm{r})}$. In addition, GRPO directly adds the KL divergence between the trained policy and the reference policy to the loss, avoiding complicating the calculation of $\hat{A}_{i,t}$. The KL divergence is estimated by the following unbiased estimator:
@@ A.2 Additional Related Work
Multimodal Chain-of-Thought. Multimodal chain-of-thought (MCoT) (Wang et al., 2025b) reasoning has recently attracted substantial attention, particularly in its integration with MLLMs. Early implementations, such as Multimodal-CoT (Zhang et al., 2024c), have established a basic MCoT pattern by generating intermediate rationales before predictions. MC-CoT (Tan et al., 2024) further refines this paradigm by employing word-level majority during training to enhance the quality of generated rationales. The dependence on high-quality MCoT training data hinders the further improvement of the inference ability of traditional methods. Most recently, the great success of Deepseek-R1 (Guo et al., 2025) has provided a way (i.e., GRPO) to enhance LLM inference capabilities through model autonomous exploration without the need for expensive CoT annotation data. Inspired by this, subsequent works utilize the GRPO strategy to efficiently enhance the reasoning ability of MLLMs. For example, Vision-R1 (Huang et al., 2025) first utilizes existing MLLM and DeepSeek-R1, as well as data filtering, through modal bridging to generate multimodal cold start CoT data, and then applies GRPO to further enhance the model’s inference capability. Perception-R1 (Yu et al., 2025a) explores the effects of RL on different perception tasks and optimizes the reward modeling to support perception policy learning. In addition, Chain-of-Shot (Hu et al., 2025) further extends GRPO strategy to optimize frame sampling via binary video summaries. In this work, we study a heatmap-based positional prior that couples MCoT with precise positional perception in a unified training framework for GRPO strategy and segmentation supervision, addressing the gap between high-level reasoning and pixel-level segmentation.
@@ B.1 Pipeline Implementation
FSDP Workers. We precompute image features offline to reduce compute, so the frozen vision backbone is excluded from the training loop. Our framework uses three FSDP workers. (i) The actor contains all trainable modules (the MLLM and the segmentation components) and is responsible for parameter updates. (ii) The rollout worker runs the MLLM only, taking image and text inputs to generate responses via next token prediction. (iii) The frozen reference worker runs an MLLM as the reference policy to compute the KL term in $\mathcal{L}_{\text{{grpo}}}$ (eq. 7) and includes the segmentation modules to decode masks used for computation of mask reward scores and group advantages.
@@
B.2 Design Details
Reward Function Design. We use a scalar mask score in $[0,1]$: given predicted mask and ground truth mask, we compute three overlap metrics (soft IoU, soft Dice, and hard IoU) and take their weighted sum with fixed coefficients $0.5$, $0.2$, and $0.3$, respectively, providing a stable localization signal for how well the prediction covers the instance. For valid outputs, the score is 1.0 by default and is reduced to 0.9 if the <think> content is longer than 2048 characters, or if any non-whitespace text appears before <think> or after the special token. Thus the five canonical cases are: invalid (0.0); valid and clean (1.0); valid but long <think> (0.9); valid but extra text before <think> (0.9); valid but extra text after the special token (0.9). For each sample, we take a weighted sum of these two components as the final reward that is assigned to the last valid response token so that GRPO updates the entire trajectory. The relative weights are specified in Section 4.4.
Algorithm 1 Generation of positional prior $\bm{H}_{\text{prior}}$
1:Image $\bm{x}_{\text{img}}$; concentration token embedding $\bm{e}_{\text{conc}}$; image encoder $\mathcal{F}_{\text{enc}}$; query head $\mathcal{F}_{\text{head}}$; fusion network $\mathcal{F}_{\text{fuse}}$; projection matrices $\{\bm{W}^{Q}_{i},\bm{W}^{K}_{i}\}_{i=1}^{n_{\text{head}}}$
2:Positional prior $\bm{H}_{\text{prior}}\in\mathbb{R}^{H\times W}$
[…]
5:for $i=1$ to $n_{\text{head}}$ do
6:  $\bm{K}_{i}\leftarrow\bm{K}\bm{W}^{K}_{i}$ $\triangleright$ $\bm{K}_{i}\in\mathbb{R}^{H\times W\times d_{h}}$
7:  $\bm{q}_{i}\leftarrow\bm{Q}\bm{W}^{Q}_{i}$ $\triangleright$ $\bm{q}_{i}\in\mathbb{R}^{d_{h}}$
8:  for $(u,v)\in\{1,\dots,H\}\times\{1,\dots,W\}$ do
9:   $S_{i}(u,v)\leftarrow\dfrac{1}{\sqrt{d_{h}}}\,\bm{q}_{i}^{\top}\bm{K}_{i}(u,v)$ $\triangleright$ $S_{i}(u,v)\in\mathbb{R}$
10:  end for
11:end for
[…]
13:return $\bm{H}_{\text{prior}}$
Positional Prior Heatmap Generation. To make the computation of the positional prior $\bm{H}_{\text{prior}}$ fully reproducible, we detail the heatmap generation procedure in Algorithm 1, starting from the image keys $\bm{K}$, the concentration query $\bm{Q}$, and the per-head scaled dot-product scores $S_{i}(u,v)$. The convolutional fusion head $\mathcal{F}_{\text{fuse}}$ then aggregates $\{\bm{S}_{i}\}_{i=1}^{n_{\text{head}}}$ into the final positional prior $\bm{H}_{\text{prior}}\in\mathbb{R}^{H\times W}$.
B.3 Training configuration
@@
B.4 Inference efficiency
Compared to representative baselines Seg-Zero and Seg-R1 using GRPO, CoPRS only adds a lightweight query head and a small extra computation for positional prior. To verify the inference efficiency of CoPRS, we conduct experiments on ReasonSeg using the same Qwen2.5-VL-3B/7B. Tables 7 and 7 report the total number of parameters, GFLOPs, and cIoU on ReasonSeg. We observe that CoPRS achieves substantially better performance under both backbones, with comparable parameter counts and inference costs.
Table 6: Comparison on 3B Models.
Method #Params(B) GFLOPs val test
Seg-R1-3B 3.97 9096.69 56.2 46.6
Seg-Zero-3B 3.97 53.1 48.6
CoPRS-3B (Ours) 4.39 9551.52 60.6 52.7
Table 7: Comparison on 7B Models.
Method #Params(B) GFLOPs val test
Seg-R1-7B 8.51 20198.96 41.2 53.7
Seg-Zero-7B 8.51 21816.71 62.0 52.0
CoPRS-7B (Ours) 8.93 22283.68 64.5 55.1
B.5 LLM usage statement
@@ Appendix C Appendix: Additional Sample Visualizations
Successful Cases. In Figure 8, we present instances from the same category and those relevant to the instruction, all showing elevated responses in the heatmap (yellow regions). More importantly, the heatmap concentrates on the instance specified by the instruction, producing a sharp peak over the target (deep red regions). This concentration guides the decoder, yielding masks with accurate boundaries. These results indicate that the MLLM reasons over the image and text input and identifies the correct referent, while the positional prior concentrates the instances for further precise mask prediction.
@@
In Figure 7, the first two rows depict scenes with many nearby instances, while the last three rows contain very small targets. Two failure modes emerge. (i) Resolution bottleneck: the positional prior is computed at 256×256 and the SAM embeddings at 64×64; when the longer image side exceeds 2k pixels, tiny objects can vanish after resizing and the decoder cannot reliably recover them. (ii) Same class crowd ambiguity: in dense groups of similar objects (e.g., crowds of people), the positional prior often spreads across many candidates with weak contrast, suggesting that a text only instruction is insufficient to disambiguate near duplicates and that the model has not fully learned the subtle semantic cues needed. These observations suggest that higher resolution inputs or multi-scale features, together with stronger instance level language grounding, are likely to improve performance on such cases.